OpenStudy (anonymous):

A used car is worth $13,500 when you buy it. The car depreciates in 8 years. Suppose you pay $3,000 down and make monthly payments of $109.38 over the 8-year period. Which set of equations best represents the line of expenses and the line of depreciation?

OpenStudy (dumbcow):

cost = 3000 + 109.38x depreciation = 13500 - (13500/96)x